How do you solve this algebra problem?
8x-2=-9+7x

Answer by checkley77(12844)

8x-2=-9+7x combine the x terms & the integers.
8x-7x=-9+2
x=-7 ans.
Proof:
8*-7-2=-9+7*-7
-56-2=-9-49
-58=-58